OpenStudy (anonymous):

Find the product: (2x^2-9y)(2x^2+9y) F= 2x^2 * 2x^2=4x^4 O= 2x^2 *9y=18x^2y I= -9y*2x^2=-18yx^2 L= -9y*9y=-81y^2 4x^4+18x^2y-18yx^2-81y^2 4x^4-81y^2 Is this right?

OpenStudy (anonymous):

The answer is right (difference of 2 squares).

OpenStudy (anonymous):

Yes, I couldn't get through that equation, could you show me how to get through that one too. so far I have. (a-b)^2=a^2 -2ab+b^2 (2x^2-9y)^2 =(2x^2)^2 -2(2x^2)(9y)+(9y)^2 2x^4-4x^2(9y)+9y^2 From there I dont know what to do. Do I *-4x^2 *9y?

OpenStudy (anonymous):

Square of A - Square of B is not the same as (A -B)^2. Your 4x^4-81y^2 is the first one and factors (A+B)(A-B).

OpenStudy (anonymous):

The answer given in the original post is correct. \[4x^4-81y^2\]